OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 5| + 2 = 5 Step 2: |x – 5| = 5 – 2 Step 3: |x – 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x – 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (anonymous):

@ParthKohli

jigglypuff314 (jigglypuff314):

once you've isolated the absolute values, you would take them out... |a-b|=c would become a-b = c and a-b = -c

OpenStudy (anonymous):

So it's C?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

Okay, Thanks! :)